# Applying Non-negative Matrix Factorization with Covariates to the Longitudinal Data as Growth Curve Model

### v1

@@ 1 Introduction
Section 2 introduces NMF with covariates and discusses the covariates composed of Gaussian kernel functions. Section 3 presents the updated formulas for optimization in the tri-NMF framework. In Section 4, we describe its relationship with GCM. Section 5 presents certain applications of NMF with covariates on real data, available on GitHub as an R package (R Core Team [19]).
@@ 2 Non-negative Matrix Factorization with covariates
Note that even if we use the same number of covariates as the number of individuals, we can obtain the optimal parameter matrix using the updated formula in Section 3. A multiple regression model with a similar covariate matrix using ridge estimators is known as the Kernel Ridge Regression Model (see Murphy [17]. Additionally, Chen et al [3] conducted a study using a general kernel-based method in NMF without covariates.
@@ 3 Optimization of base matrix and parameter matrix
We measure the discrepancy between $Y$ and $\hat{Y}$ by 1) the squared Euclidean distance and 2) Kullback-Leibler (KL) divergence. Minimizing each objective function is equivalent to applying the maximum likelihood method when the observation is distributed as 1) Normal, and 2) Poisson. The updated formulas for tri-NMF with a squared Euclidean distance were also presented by Ding et al [6] and Čopar et al [5] when $X$ and $A$ are orthogonal matrices. For the numerical computation of NMF without covariates, the package NMF of the statistical software R (R Core Team [19]) by Gaujoux and Seoighe [8] is available.
@@ 3.1 Squared Euclidean distance
When the coefficient matrices $B=\Theta A$ and $\hat{Y}$ are given, the updated formula for the basis matrix $X$ remains the same as that without using covariates.
@@
Next, we present the updated formula for the parameter matrix $\Theta$ when the basis matrices $X$ and $\hat{Y}$ are as follows:
@@ 5.1 Growth curve data on an orthodontic measurement
> library(nlme)
> Y <- matrix(Orthodont\$distance,nrow=4)
@@
> library(nmfkc)
> result <- nmfkc(Y,Q=2,epsilon=1e-8)
@@ 5.2 Daily temperature in Canada
> library(fda)
> data(CanadianWeather)
> d <- CanadianWeather\$dailyAv[,,1]
> Y <- d-min(d)
> library(nmfkc)
> result <- nmfkc(Y,Q=2)
> result\$r.squared
[1] 0.9853615
@@
Fig.4 shows (a) the column vectors of the basis matrix, and (b) the probabilities at the weather stations. The sum of each basis vector is 1. The fitted value for each weather station was approximated as the sum of the coefficient multiples of the two bases. The two coefficients are non-negative and can be interpreted as probabilities, as shown in the star chart with normalized longitudes and latitudes. As shown in Fig.4(a), basis 1 represents the temperature variation in the Northern Hemisphere, with lower temperatures in winter and higher temperatures in summer. Basis 2, conversely, is shaped to push up winter temperatures. In Fig.4(b), the probability of basis 1 is higher north of the interior, and the probability of basis 2 increases as basis 1 moves southward. Additionally, the probability of basis 2 tends to be higher on the west coast than on the east coast, even in the same southern area. These trends suggest that the coefficient matrix can be explained by location information: longitude and latitude.
@@
> result <- nmfkc(Y,A,Q=2)
> result\$r.squared
[1] 0.9340371
@@
Finally, we attempted to use the basis matrix obtained in the case of the growth curve model, but the determinant of $S$ was zero and the inverse matrix could not be obtained; thus, $\hat{\Theta}$ in (16) could not be computed. Similarly, when the Gaussian kernel function is used as a covariate, the eigenvalue of $AA^{\prime}$ is 0, and the inverse matrix cannot be computed, resulting in a failure to obtain $\hat{\Theta}$. Thus, GCM may be difficult to apply when the dimension $P$ of the observed vector is high.
@@ 5.3 US presidential inaugural address texts
We treated the inaugural speeches of U.S. presidents as textual data observed over time. The text data were stored as data_corpus_inaugural in the quanteda package and contained 59 speeches from President Washington in 1789 to President Biden in 2021. In Script 6, we first extracted from the text data 40 frequent words with 5 or more characters and 100 or more occurrences. Generally, a frequency table in which the rows are words, and the columns are documents is called a term–document matrix. If we apply NMF to the term document matrix, then such can be interpreted as a Topic Model. In other words, each column vector of the basis matrix is the probability of a word constituting a topic, and the proportion of coefficients can be interpreted as topic probability. The rows of the basis matrix are all probabilities of the same word for each topic, and the proportion of probabilities provides the magnitude of the word’s contribution to each topic.
Script 6: NMF as Topic Model
@@
> library(nmfkc)
> result <- nmfkc(Y,Q=3)
> result\$r.squared
[1] 0.6562806
@@
> result <- nmfkc(Y,A,Q=3)
> result\$r.squared
[1] 0.3279501
@@
NMF was then applied by creating a covariate matrix $A$ by using the Gaussian kernel function from the year of inauguration. The coefficient of determination was 0.328, which was worse. Bar chart of topic probabilities for each president obtained under the model are shown in Fig.6(b). The general trend is similar to that shown in Fig.6(a), but the Gaussian kernel function smoothes out the vertical variation seen in Fig.6(a), making Fig.6(b) more straightforward to interpret. Note that when $X$ optimized by NMF with covariates is used for GCM to find $\hat{\Theta}$ in (16), $S$ cannot be calculated because the determinant of $AA^{\prime}$ is 0.
@@ 5.4 Bone mineral density in adolescents
In Script 7, the dimension of the observation matrix $Y$ is $P=1$, resulting in a scalar basis matrix with $X=1$ and $Q=1$. Hence, the NMF with covariates is equivalent to a multiple regression model that can be written as the updated equation (18). The Gaussian kernel function for age and the function of the product of age and male dummy variables were used to obtain smoothed curves for gender, see Fig.7. Since the covariate matrix has $R=970$ rows and $N=485$ columns, there are $R$ parameters to be optimized that are greater than the number of individuals $N$. Therefore, because this model is formally included in a multiple regression model, the least-squares estimator can be written as $\hat{{\bm{\theta}}}=(AA^{\prime})^{-1}A{\bm{y}}$. However, it cannot be calculated because the determinant of $AA^{\prime}$ is 0. For example, ridge estimators might be required.
Script 7: NMF with the covariates as Multiple Regression Model
@@
> d <- bone
> d\$male <- ifelse(d\$gender=="male",1,0)
@@
> result <- nmfkc(Y,A=rbind(A1,A2),Q=1)
> result\$r.squared
[1] 0.4264267

### v3

@@ 1 Introduction
Section 2 introduces NMF with covariates and discusses the covariates composed of Gaussian kernel functions. Section 3 presents the update formulas for optimization in the tri-NMF framework. In Section 4, we describe its relationship with GCM. Section 5 presents certain applications of NMF with covariates on real data, available on GitHub as an R package (R Core Team [19]).
@@ 2 Non-negative Matrix Factorization with covariates
Note that even if we use the same number of covariates as the number of individuals, we can obtain the optimal parameter matrix using the update formula in Section 3. A multiple regression model with a similar covariate matrix using ridge estimators is known as the Kernel Ridge Regression Model (see Murphy [17]. Additionally, Chen et al [3] conducted a study using a general kernel-based method in NMF without covariates.
@@ 3 Optimization of base matrix and parameter matrix
We measure the discrepancy between $Y$ and $\hat{Y}$ by 1) the squared Euclidean distance and 2) Kullback-Leibler (KL) divergence. Minimizing each objective function is equivalent to applying the maximum likelihood method when the observation is distributed as 1) Normal, and 2) Poisson. The update formulas for tri-NMF with a squared Euclidean distance were also presented by Ding et al [6] and Čopar et al [5] when $X$ and $A$ are orthogonal matrices. For the numerical computation of NMF without covariates, the package NMF of the statistical software R (R Core Team [19]) by Gaujoux and Seoighe [8] is available.
@@ 3.1 Squared Euclidean distance
When the coefficient matrices $B=\Theta A$ and $\hat{Y}$ are given, the update formula for the basis matrix $X$ remains the same as that without using covariates.
@@
Next, we present the update formula for the parameter matrix $\Theta$ when the basis matrices $X$ and $\hat{Y}$ are as follows:
@@ 5 Examples of data analysis
Four datasets are treated here, but ordinary GCM was only applied to the first 4.1. For the other three, the maximum likelihood estimator could not be obtained in GCM: GCM is suitable for making statistical inferences, but was difficult to apply to observations of various sizes.
@@ 5.1 Growth curve data on an orthodontic measurement
> library(nlme); Y <- matrix(Orthodont\$distance,nrow=4)
@@
> library(nmfkc); result <- nmfkc(Y,Q=2,epsilon=1e-8)
@@ 5.2 Daily temperature in Canada
> library(fda); data(CanadianWeather)
> d <- CanadianWeather\$dailyAv[,,1]; Y <- d-min(d)
> library(nmfkc); result <- nmfkc(Y,Q=2); result\$r.squared
[1] 0.9853628
@@
Fig.4 shows (a) the column vectors of the basis matrix, and (b) the probabilities at the weather stations. The sum of each basis vector is 1. The fitted value for each weather station was approximated as the sum of the coefficient multiples of the two bases. The two coefficients are non-negative and can be interpreted as probabilities, as shown in the star chart with normalized longitudes and latitudes. As shown in Fig.4(a), basis 1 represents the temperature variation in the Northern Hemisphere, with lower temperatures in winter and higher temperatures in summer. Basis 2, conversely, is shaped to push up winter temperatures. In Fig.4(b), the probability of basis 1 is higher north of the interior, and the probability of basis 2 increases southward. Additionally, the probability of basis 2 tends to be higher on the west coast than on the east coast, even in the same southern area. These trends suggest that the coefficient matrix can be explained by location information: longitude and latitude.
@@
> library(nmfkc); result <- nmfkc(Y,A,Q=2); result\$r.squared
[1] 0.9340437
@@
From equation (3), the probability of basis 2 can be derived on new locations and it is obtained in B.prob of Script 6 and those can be visualized as Fig.6. The general trend is the same as in Fig.4(b).
Script 6: Soft clustering on new covariates
> v <- seq(from=0,to=1,length=20)
> V <- t(cbind(expand.grid(v,v)))
> A <- nmfkc.kernel(U,V,beta=best.beta);
> B <- result\$C %*A; B.prob <- prop.table(B,2)
Figure 6: Geographic distribution of the probability of basis 2. The contours lines shows the probability of basis 2 and those are obtained as smooth curves based on Gaussian kernel functions.
@@ 5.3 US presidential inaugural address texts
We treated the inaugural speeches of U.S. presidents as textual data observed over time. The text data were stored as data_corpus_inaugural in the quanteda package and contained 59 speeches from President Washington in 1789 to President Biden in 2021. In Script 7, we first extracted from the text data 40 frequent words with 5 or more characters and 100 or more occurrences. Generally, a frequency table in which the rows are words, and the columns are documents is called a term–document matrix. If we apply NMF to the term document matrix, then such can be interpreted as a Topic Model. In other words, each column vector of the basis matrix is the probability of a word constituting a topic, and the proportion of coefficients can be interpreted as topic probability. The rows of the basis matrix are all probabilities of the same word for each topic, and the proportion of probabilities provides the magnitude of the word’s contribution to each topic.
Script 7: NMF as Topic Model
@@
> library(nmfkc); result <- nmfkc(Y,Q=3); result\$r.squared
[1] 0.6563144
@@
> result <- nmfkc(Y,A,Q=3); result\$r.squared
[1] 0.3280116
@@
NMF was then applied by creating a covariate matrix $A$ by using the Gaussian kernel function from the year of inauguration. The coefficient of determination was 0.328, which was worse. Bar chart of topic probabilities for each president obtained under the model are shown in Fig.7(b). The general trend is similar to that shown in Fig.7(a), but the Gaussian kernel function smoothes out the vertical variation seen in Fig.7(a), making Fig.7(b) more straightforward to interpret.
@@ 5.4 Bone mineral density in adolescents
In Script 8, the dimension of the observation matrix $Y$ is $P=1$, resulting in a scalar basis matrix with $X=1$ and $Q=1$. Hence, the NMF with covariates is equivalent to a multiple regression model that can be written as the update equation (18). The Gaussian kernel function for age and the function of the product of age and male dummy variables were used to obtain smoothed curves for gender, see Fig.8.
Script 8: NMF with the covariates as Multiple Regression Model
@@
> d <- bone; d\$male <- ifelse(d\$gender=="male",1,0)
@@
> result <- nmfkc(Y,A=rbind(A1,A2),Q=1); result\$r.squared
[1] 0.4264902